"Solve \t\u222b\u2592\u221bx dx"


1
Expert's answer
2022-06-08T04:57:56-0400

"\int {\sqrt[3]{x}} dx = \int {{x^{\frac{1}{3}}}} dx = \frac{{{x^{\frac{1}{3} + 1}}}}{{\frac{1}{3} + 1}} + C = \frac{{{x^{\frac{4}{3}}}}}{{\frac{4}{3}}} + C = \frac{3}{4}x\sqrt[3]{x} + C"

Answer: "\frac{3}{4}x\sqrt[3]{x} + C"